OpenStudy (anonymous):

hello, i do have te answers but i'd like to make sure i have this right before graphing the circle. and thank you in advance for your help. *a circle has the equation 3(x-4)^2+3y^2=27 Find the center (h,k) and radius r. *Then find the intercepts, if any, of the graph.

hero (hero):

(h,k) = (4,0)

hero (hero):

r = sqrt{27} or 3sqrt{3}

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

Need to bring the 3 out. \[r = \sqrt{27/3}\] r =3

OpenStudy (anonymous):

Hello Hero and Nilanshki, yes, I have the same exact answer and had plotted the circle but I was worried I was worried. I'm new to this forum...glad we can all help each other here. Thanks so much for helping me by confirming it.

OpenStudy (anonymous):

I meant I was worried about not getting the right answer. lol. thnaks again.

OpenStudy (anonymous):

"Need to bring the 3 out." By which i mean 3(x-4)^2+3y^2=27 3((x-4)^2+y^2)=27 (x-4)^2+y^2= 9 (27/3)

OpenStudy (anonymous):

Hi Dangolbery, SOunds great. I actually have the answers written down before asking...just had to make sure. Thanks so much for your help too.

OpenStudy (anonymous):

i'm doing the problems by hand cause i can't figure out y new calculator yet. glad we have this site. you guys are awesome!

OpenStudy (nilankshi):

good